Using Python 3+ for Question P5.9

Instructions: Use one main() to call each of the functions you created. Only use one value of r and h for all the function. Ask the user for the r and h in the main() as an input, and h for all the functions. You should put the functions for areas in a separate file. ​For example, firstDigtec7ze Write a function e digits n) (returning the number of digits in the argument returning the last digit of the argument) is 1,1astDec2) is 9, and igitsC12) is 4 def middle(string) that returns a string containing t odd, or the two middle characters if returns dd a string containing the midle in le character in string if the length of string is is even. For example, middeCnide) P5.5 Write a function def repeat(string, n, delim) that returns the string string repeated n times, separated by the string delte. For example, repeat(ho, 3, . ) returns ho, ho, ho. P5.6 Write a function def countVowe1s(string) that returns a count of all vowels in the string string. Vowels are the letters a, e,i,o, and u, and their uppercase variants .. P5.7 Write a function def countWords(string) that returns a count of all words in the string string. Words are separated by spaces. For example, countwords(Mary had a little 1amb) should return 5. It is a well-known pheno words have two characters flipped, provided the first and last letter of each word are not changed. For example, P5.8 menon that most people are easily able to read a text whose I dnot gvie a dman for a man taht can olny seplla wrod one way. (Mrak Taiwn) Write a function scramble (word) that constructs a scrambled version of a given word, randomly flipping two characters other than the first and last one.Then write a program that reads words and prints the scrambled words. Write functions def sphereVolume(r) def sphereSurface(r) def cylinderVolume (r, h) def cylinderSurface(r, h) def coneVolume(r, h) def coneSurface(r, h) that compute the volume and surface area of a sphere with radius r, a cylinder with a circular base with radius r and height h, and a cone with a circular base with radius r and height h. Then write a program that prompts the user for the values of r and h, calls the six functions, and prints the results. P5.10 Write a function def readFloat (prompt)
Answer #1

from math import pi



def sphereSurface(r):
    area1=(4*pi*r**2)
    return area1

def sphereVolume(r):
    volume1=((4/3)*pi*r)**3
    return volume1

def cylinderSurface(r,h):
    area2=(2*pi*r**2+2*pi*r*h)
    return area2

def cylinderVolume(r,h):
    volume2=(pi*r**2*h)
    return volume2

def coneSurface(r,h):
    area3=(pi*r+pi*r**2)
    return area3

def coneVolume(r,h):
    volume3=((1/3)*pi*r**2*h)
    return volume3


r = float(input('Enter radius:'))
h = float(input('Enter height:'))

print('Spehere Volume:', sphereVolume(r))
print('Spehere Surface:', sphereSurface(r))
print('Cylinder Volume:', cylinderVolume(r, h))
print('Cylinder Surface:', cylinderSurface(r, h))
print('Cone Volume:', coneVolume(r, h))
print('Cone Surface:', coneSurface(r, h))

Python 3.6.1 (default, Dec 2015, 13:05:11) [GCC 4.8.2] on linux saved main.py Enter radius: 3.4 Enter height: 2.3 Spehere VolHi. please find the answer.. In case of any doubts, please ask in comments. If the answer helps you, please upvote. Thanks!